Please Answer within the hour and Show work. Thank you.No answer submission is required for this problem during this quiz. Please read the instructions and then go to the next prob
Answer #1

A) 1) Lewis structure for all three resonance :NEN -Ö: N=N=ö <>. nzö: (11) (11) 2. Formal charge on each atom in each resorenF C structure TiT on 0 6-C1246+2). + on middle al = S-48eati on terminal N = 5-(2x246) = -2 FC F.C +) No ON Only two more decActual Bond lenght of N-0 is 19pm. in M20, which lies blw N-o single bonel lenght (147pm) and unght clispm). This can be possplease rate if you find helpfull

Thank you